FTIR SPECTRUM AND PERTURBATION ANALYSIS OF THE ν_{2} BAND OF ^{15}NO_{3}

N. SHIMIZU, R. FUJIMORI, J. TANG, \underline{K. KAWAGUCHI}, Department of Chemistry, Faculty of Science, Okayama University, 3-1-1, Tsushima-Naka, Okayama 700-8530, Japan; T. ISHIWATA, Graduate School of Information Sciences, Hiroshima City University, 3-4-1 Otsuka-Higashi, Hiroshima 731-3194, Japan,.

In 1987, Friedl and Sander^{a} reported observation of the FTIR spectrum of the ν_{2} band of ^{14}NO_{3} in 762 cm^{−1} region, and rotational analysis was carried out for only transitions with K=3. The obtained negative centrifugal distortion constant D_{N} implied presence of perturbation from other states, but the perturbation analysis has not been carried out. In the present study, the ν_{2} band of ^{15}NO_{3} was first observed in 742 cm^{−1} region. In contrast to the case of ^{14}NO_{3} , rotational assignments were carried out for K=0, 3, 6, 9, 12, and 15, and transitions with N≤12 were fitted with usual energy level expression for D_{3h} molecule. However, when we included higher N transitions up to N=22, an effect of Coriolis interaction from v_{4}=2 appeared in negative D_{N} for K=0 and staggering in K=3. An energy matrix including ℓ=0 and ℓ=±2 of v_{4}=2 and ℓ=0 of v_{2} =1 was used to obtain the energy values of v_{2} =1 , including the Corilis and ℓ-type resonance terms. From the determined interaction constant between v_{2} =1 and v_{4}=2 , anharmonic constant k_{444} was estimated, by assuming a mixing of v_{4}=1 and v_{4}=2.

aR. R. Friedel and S. P. Sander, J. Phys. Chem. 91, 2721 (1987).